I need help to solve this problem seems I can get the right answer
Solve
3(x – 2) + 7x = 8x + 0

Answer by checkley75(3666)

3(X-2)+7X=8X+0
3X-6+7X=8X
3X+7X-8X=6
2X=6
X=6/2
X=3 ANSWER.
PROOF
3(3-2)+7*3=8*3
3*1+21=24
3+21=24
24=24